OpenStudy (anonymous):

The total charge of a nucleus = 92*1.6*10^-19 Mass of Nucleus = A * 1.67 * 10^-27 Rearrange to find A, then subtract the number of protons, the answer should be about 146. Can someone help me find A?

OpenStudy (anonymous):

A =146 but how do i get to that?

OpenStudy (anonymous):

The specific charge is 3.7*10^7

OpenStudy (anonymous):

hey is the value of A 188? bcoz,the first statement says that there are 94 protons ....... in the next statement,A=protons+neutrons in the nucleus no of protons=no of neutrons so, A=94+94=188.

OpenStudy (anonymous):

No A = 146. I am looking to find out how i arrive at that number. Sorry there are 92 protons

OpenStudy (anonymous):

lol.yeah 92!

OpenStudy (anonymous):

specific charge=charge/mass so \[(92*1.6*10^-19)/(A*1.67*10^-27) =3.7*10^7\] calculate and u get A=238.23 A-92=146.23

OpenStudy (anonymous):

which standard are u?

OpenStudy (anonymous):

I am only Gcse

OpenStudy (anonymous):

Gcse means?

OpenStudy (anonymous):

Below A-level?

OpenStudy (anonymous):

just calculate or use a scientific calculator!!!!

OpenStudy (anonymous):

I don,t think you understand what i am saying. You have given me the answer 3.7*10^7. That is ok. But how do you get the answer 238?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The specific charge of 3.7*10^7 is for Uranium 238?